# Panoramic Out-of-Distribution Segmentation for Autonomous Driving

### v2

@@ VI Conclusion
In the future, we intend to incorporate multimodal fusion techniques, combining various scene information types (e.g., depth and point clouds) to boost the model’s adaptability to defocus and motion blur. Additionally, incorporating time series information to strengthen the scene perception capabilities in dynamic environments is worth investigating to further improve PanOoS performance, with the aim of adapting to complex dynamic conditions, as well as open-set and 3D perception tasks.
@@ A-A Prompt Templates
We adopt the prompt templates from [31, 30], but make additional prompt optimization to describe the different distribution items more accurately. Specifically, the lists of templates for class and distribution prompts are as follows:
Class:
@@ B-A1 DenseOoS
As shown in Table B.1, we present the specific types of outliers in DenseOoS. DenseOoS is based on the designs in POC [81] and additionally includes $5$ other common anomalous objects (cattle, duck, cock, box, and rock). To ensure the high quality and rationality of the generated images, particularly that the outliers are generated in realistic areas within the real street backgrounds, our screening process follows the “preliminary screening $\rightarrow$ independent screening $\rightarrow$ voting” workflow.
The overall process of dataset generation and construction is as follows. First, all text prompts for image inpainting (a good photo of a $\{class\}$) are determined based on the DensePASS validation dataset; Next, we select appropriate regions to be inserted (e.g., the road and sidewalk) and the selected rectangular area is scaled to match the resolution of the real panoramic images. Then, we set different random seeds and use the POC algorithm to batch-generate images and outlier annotations. Finally, the suitable images for constructing the DenseOoS dataset are determined through preliminary manual screening, independent screening, and majority voting.
B-A2 QuadOoS
In addition, we also list some types of outliers in QuadOoS. As present in Fig. 16, the outliers mainly consist of two parts: human-defined and existing in the neighborhood.
@@ B-A2 QuadOoS
B-B Comparative Experiments
For the experiments on PanOoS, we compare POS to the state-of-the-art pinhole-OoS methods: per-pixel architecture [32, 33] and mask-transformer-based [26, 28, 27] methods, using the models that were shared in their respective repositories.
@@ C-A More in-distribution segmentation results
DAFormer [34] 48.94 31.78 84.12 26.51 31.06 24.92 6.83 7.73 71.86 40.63 91.23 44.13 2.1 78.02 55.02 25.24 84.89 61.37 29.59 44.53
Trans4PASS (S) [12] 74.28 30.68 84.79 28.45 36.60 27.90 15.40 15.11 78.99 33.71 93.28 49.69 6.79 82.00 46.42 51.79 72.84 68.34 40.63 49.35
Trans4PASS+ (S) [2] 74.74 34.17 85.43 23.46 42.78 28.45 17.14 18.08 79.64 29.41 93.00 48.74 5.40 82.76 50.54 44.91 91.49 61.49 40.20 50.11
360SFUDA++ [48] 72.29 45.24 84.91 32.11 38.18 29.83 19.28 17.98 76.02 27.23 92.01 53.14 18.74 81.35 44.42 45.38 82.84 60.93 36.92 50.46
@@
To provide a thorough analysis of the panoramic segmentation and demonstrate that excellent closed-set segmentation performance ensures effective out-of-distribution segmentation, we present the accuracy results for each inlier class in the DenseOoS dataset, as shown in Table C.1. To benchmark our model, we conduct a comparative study with the previous state-of-the-art panoramic semantic segmentation methods, namely DAFormer [34], Trans4PASS [12], Trans4PASS+ [2], and 360SFUDA++ [48].
DAFormer introduces domain-adaptive transformers to improve semantic segmentation across domains. Trans4PASS and its enhanced variant, Trans4PASS+, employ transformer-based architectures specifically designed to mitigate distortions and exploit contextual dependencies in panorama. 360SFUDA++ represents an advanced source-free domain adaptation framework based on self-training, tailored to panoramic scenarios. While these methods achieve good performance under closed-set conditions, they experience notable degradation in the presence of out-of-distribution regions, owing to their limited generalization beyond the predefined semantic space. This limitation underscores the importance of enhancing robustness to unexpected content in real-world panoramic scenes.
As shown in the experimental results, our model achieves a significant mIoU of $63.30\%$, surpassing the performance of the previous best model by a large margin of $12.84\%$. A detailed breakdown of per-class performance reveals that our model excels across numerous categories, including road, sidewalk, building, wall, fence, vegetation, and terrain etc. The excellent performance in these classes demonstrates that our model is able to adapt well to the background clutter and pixel distortions introduced by the broader FoV of panoramic images, thus achieving leading segmentation performance.
@@ C-B Additional Ablation Study
Features and layers in Transformer Decoder. To further analyze the most favorable settings for PanOoS, we conduct additional ablation experiments on the decoded features and the corresponding number of layers in the transformer decoder. As shown in Table C.2, the results align with previous work: by decoding only semantically rich features and reducing the number of decoder layers, semantic confusion can be effectively alleviated, which helps to better align the logit scores and results in better outlier performance [27]. Specialized object queries perform worse with more decoder layers. This also supports our view that, while ensuring semantic clarity, enhancing the model’s perception of panoramic pixels can significantly improve the performance of PanOoS.
@@ Appendix D More Qualitative Results
We showcase more visualization results of panoramic out-of-distribution segmentation of POS on DenseOoS and QuadOoS in Figs. D.1 and D.2. These examples demonstrate that POS achieves excellent performance on the PanOoS benchmark. Compared to other mask-transformer-based methods: EAM [26], Mask2Anomaly [28], and RbA [27], POS reduces false positives on the ambiguous background regions and boundaries of inliers while remaining sensitive to outliers. These improvements can be observed more prominently on QuadOoS (see Fig. D.2) under defocus and motion blur scenarios.
@@ E-A Societal impacts
In this study, we have introduced a novel task called Panoramic Out-of-distribution Segmentation (PanOoS) and established a comprehensive benchmark incorporating various well-known baseline models. We found that these baseline models exhibit limited performance on the PanOoS task, primarily due to the background clutter and pixel distortions introduced by the panoramic broader FoV. To address this issue, we have developed POS, a solution that significantly enhances the performance of the PanOoS benchmark, outperforming existing pinhole-OoS methods: pixel-wise architecture and mask-transformer-based method, and achieving promising state-of-the-art results.

### v3

@@ A-A Prompt Templates
We adopt the prompt templates from [radford2021learning, yu2023convolutions], but further optimize them to more accurately describe items from different distributions. Specifically, the class and distribution prompt templates are as follows:
A-A1 Class
@@ B-A1 DenseOoS
Figure B.1: Examples from the DenseOoS dataset. Image backgrounds are highly cluttered, and distortions are severe. Outliers appear in semantically plausible regions. Zoom in for a better view.
The specific types of outliers in DenseOoS are listed in Table B.1. The normal labels cover $19$ classes that align with those of Cityscapes [cordts2016cityscapes], while the abnormal labels contain $30$ classes of outliers, including animals, garbage, and furniture. DenseOoS additionally includes $5$ other common anomalous objects (cattle, duck, cock, box, and rock) compared to POC [de2024placing]. To ensure the high quality and rationality of the generated images, particularly that the outliers are generated in realistic areas within the real street scenes, our screening process follows the “preliminary screening $\rightarrow$} independent screening $\rightarrow$} voting” workflow, eliminating unreasonable object locations and normal situations, e.g., birds flying in the city sky or garbage/garbage bags/wheelie bin placed on the sidewalk.
The overall process of dataset generation and construction is as follows. First, all text prompts for image inpainting (a good photo of a $\{class\}$) are determined based on the DensePASS validation dataset. Next, we select appropriate regions to be inserted (e.g., the road and sidewalk) and the selected rectangular area is scaled to match the resolution of the real panoramic images. Then, we set different random seeds and use the POC algorithm to batch-generate images and outlier annotations. Finally, the suitable images for constructing the DenseOoS dataset are determined through preliminary manual screening, independent screening, and majority voting. Specifically, a total of $14,031$ images are generated, of which $1,244$ are retained after preliminary manual screening. Following an independent screening and voting process by three examiners, we ultimately select $1,000$ images for the construction of DenseOoS. As shown in Fig. B.1, the outliers are diverse and distributed within semantically plausible regions of the images. A significant portion of these outliers extends beyond the pinhole narrow FoV, making them particularly suitable for the evaluation of the PanOoS task.
B-A2 QuadOoS
In addition, we list some types of outliers in QuadOoS. As presented in Fig. B.2, the outliers mainly consist of two parts: human-defined objects and outliers existing in the neighborhood. Compared to panoramic images captured under static conditions, the images in QuadOoS (see Fig. B.3) pose significant challenges to existing methods, such as uneven exposure, color inconsistencies due to the broader FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect. Each image contains an average of $7$ outliers, accurately reflecting real-world situations and posing substantial challenges for PanOoS methods.
@@ B-A2 QuadOoS
Figure B.3: Examples from the established QuadOoS dataset. The images come with challenges, including uneven exposure, color inconsistencies due to the broader FoV, and increased motion blur. Numerous small outliers are scattered across the scene.
B-B Evaluation Metrics
In the context of the Panoramic Out-of-distribution Segmentation (PanOoS) benchmark, we employ three metrics, namely Area under PRC Curve (AuPRC), False Positive Rate at a true positive rate of $95\%$ (FPR95), and mean Intersection-over-Union (mIoU), to evaluate the performance of models. For pixel-wise evaluation, $\mathcal{G}{\in}\{\mathcal{G}_{in},\mathcal{G}_{out}\}$ is the annotated ground truth labels for an image containing outliers, where $\mathcal{G}_{in}$ and $\mathcal{G}_{out}$ represent the inlier and outlier labels, respectively. Let $\mathcal{G}_{p}$ denote the model prediction obtained by a threshold $\tau$. The True Positive Rate (TPR), precision, and recall can be written as:
The AuPRC can be approximated as:
$AuPRC\!$ $=\!\int_{\tau}\!precision(\tau)\,recall(\tau).$ (28)
where $\tau_{95}$ is the threshold at which the TPR equals 95%.
$mIoU\!=\frac{1}{K}\!\sum_{k=1}^{K}\!\frac{p_{k}\!\cap\!g_{k}}{p_{k}\!\cup\!g_{k}}.$ (30)
B-C Comparative Experiments
For the experiments on PanOoS, we compare POS to the state-of-the-art pinhole-OoS methods: per-pixel architectures [tian2022pixel, grcic2022densehybrid] and mask-transformer-based [grcic2023advantages, rai2024mask2anomaly, nayal2023rba] methods, using the models that were shared in their respective repositories.
@@ C-A More in-distribution segmentation results
DAFormer [hoyer2022daformer] 48.94 31.78 84.12 26.51 31.06 24.92 6.83 7.73 71.86 40.63 91.23 44.13 2.1 78.02 55.02 25.24 84.89 61.37 29.59 44.53
Trans4PASS (S) [zhang2022bending] 74.28 30.68 84.79 28.45 36.60 27.90 15.40 15.11 78.99 33.71 93.28 49.69 6.79 82.00 46.42 51.79 72.84 68.34 40.63 49.35
Trans4PASS+ (S) [zhang2024behind] 74.74 34.17 85.43 23.46 42.78 28.45 17.14 18.08 79.64 29.41 93.00 48.74 5.40 82.76 50.54 44.91 91.49 61.49 40.20 50.11
360SFUDA++ [zheng2024360sfuda++] 72.29 45.24 84.91 32.11 38.18 29.83 19.28 17.98 76.02 27.23 92.01 53.14 18.74 81.35 44.42 45.38 82.84 60.93 36.92 50.46
@@
To provide a thorough analysis of panoramic segmentation and demonstrate that strong closed-set segmentation performance ensures effective out-of-distribution segmentation, we present the accuracy results for each inlier class in the DenseOoS dataset, as shown in Table C.1. To benchmark our model, we conduct a comparative study with the existing state-of-the-art panoramic semantic segmentation methods, namely DAFormer [hoyer2022daformer], Trans4PASS [zhang2022bending], Trans4PASS+ [zhang2024behind], and 360SFUDA++ [zheng2024360sfuda++]. DAFormer introduces domain-adaptive transformers to improve semantic segmentation across domains. Trans4PASS and its enhanced variant, Trans4PASS+, employ transformer-based architectures specifically designed to mitigate distortions and exploit contextual dependencies in panoramas. 360SFUDA++ represents an advanced source-free domain adaptation framework based on self-training, tailored to panoramic scenarios. While these methods achieve good performance under closed-set conditions, they experience notable degradation in the presence of out-of-distribution regions, owing to their limited generalization beyond the predefined semantic space. This limitation underscores the importance of enhancing robustness to unexpected content in real-world panoramic scenes.
As shown in the experimental results, our model achieves a significant mIoU of $63.30\%$, surpassing the performance of the previous best model by a large margin of $12.84\%$. A detailed breakdown of per-class performance reveals that our model excels across numerous categories, including road, sidewalk, building, wall, fence, vegetation, and terrain etc. The excellent performance in these classes demonstrates that our model is able to adapt well to the pixel distortions and background clutter introduced by the broader FoV of panoramic images, thus achieving leading segmentation performance.
@@ C-B Additional Ablation Study
Num Iter AuPRC FPR95 mIoU
15kk 79.19 0.49 57.49
20kk 78.22 0.57 60.14
25kk 84.04 0.51 62.71
30kk 85.56 0.45 63.30
35kk 82.08 0.56 63.14
((a))
$p_{out}$ AuPRC FPR95 mIoU
0.1 86.30 0.45 62.47
0.2 85.72 0.45 62.73
0.3 85.56 0.45 63.30
0.4 81.28 0.59 62.17
0.5 80.97 0.53 62.36
((b))
Module AuPRC FPR95 mIoU
w/o 55.66 7.44 54.11
Decoders 84.25 0.50 62.38
Pixel Dec. 85.56 0.45 63.30
Trans. Dec. 46.91 5.58 57.89
((c))
TABLE C.2: Ablation study on outlier exposure fine-tuning.
C-B1 Outlier Exposure Fine-tuning
We further empirically evaluate the impact of fine-tuning parameters during the optimization of the prompt distribution space. As shown in Table LABEL:tab:ablationIter, results on the DenseOoS dataset demonstrate that the model reaches optimal performance after approximately $30k$ iterations, after which overfitting begins. For the outlier data exposure control parameter $p_{out}$, experiments reveal that when $p_{out}{>}0.3$, both closed-set performance and out-of-distribution segmentation performance deteriorate. Based on these results, we selected $0.3$ as the optimal value, balancing outlier detection with closed-set performance.
@@ C-B1 Outlier Exposure Fine-tuning
C-B2 Features and layers in Transformer Decoder
To further analyze the most favorable settings for PanOoS, we conduct additional ablation experiments on the decoded features and the corresponding number of layers in the transformer decoder. As shown in Table C.3, the results align with previous work: decoding only semantically rich features and reducing the number of decoder layers effectively alleviates semantic confusion, which helps to better align the logit scores and results in better outlier performance [nayal2023rba]. Specialized object queries perform worse with more decoder layers. This also supports our view that, while ensuring semantic clarity, enhancing the model’s perception of panoramic pixels can significantly improve the performance of PanOoS.
@@ Appendix D More Qualitative Results
In Fig. C.1, we visualize the PanOoS results produced by POS and other mask-transformer-based methods. Notably, POS effectively suppresses false positives in background regions, especially at the boundaries separating inlier regions, while better preserving the smoothness of the outlier map compared to other methods, despite sharing the same mask classification training paradigm.
Figure D.0: Qualitative comparisons of PanOoS. When existing pinhole-OoS methods are transferred to PanOoS, they tend to misidentify complex background regions as outliers. Although EAM [grcic2023advantages] has a low false positive rate, it incorrectly identifies bicycles as outliers and nearly overlooks the actual outliers. In contrast, Mask2Anomaly [rai2024mask2anomaly] and RbA [nayal2023rba] suffer from significant false positives in the background regions, whereas POS delivers more precise PanOoS across the $360^{\circ}$.
We showcase more visualization results of panoramic out-of-distribution segmentation of POS on DenseOoS and QuadOoS in Figs. D.1 and D.2. These examples demonstrate that POS achieves excellent performance on the PanOoS benchmark. Compared to other mask-transformer-based methods, EAM [grcic2023advantages], Mask2Anomaly [rai2024mask2anomaly], and RbA [nayal2023rba], POS reduces false positives on the ambiguous background regions and boundaries of inliers while remaining sensitive to outliers. These improvements can be observed more prominently on QuadOoS (see Fig. D.2) under defocus and motion blur scenarios.
@@ E-A Societal impacts
In this study, we have introduced a novel task called Panoramic Out-of-distribution Segmentation (PanOoS) and established a comprehensive benchmark incorporating various well-known baseline models. We found that these baseline models exhibit limited performance on the PanOoS task, primarily due to the pixel distortions and background clutter introduced by the panoramic broader FoV. To address this issue, we have developed POS, a solution that significantly enhances the performance of the PanOoS benchmark, outperforming existing pinhole-OoS methods: pixel-wise architectures and mask-transformer-based methods, and achieving state-of-the-art results.
@@
E-B Future work
Although POS aims to deal with pixel distortions and background clutter in panoramic images and shows strong performance, we note that the model is not ideal for defocus and motion blur adaptation in intense motion scenarios. These issues result in performance degradation in complex dynamic environments. To address these limitations, future studies could integrate multi-modal fusion approaches to combine various types of scene information, e.g. depth and point clouds, to enhance the model’s adaptability to defocus and motion blur. Additionally, incorporating temporal information to enhance the scene perception capabilities in dynamic environments is worth investigating to further enhance PanOoS performance, with the aim of adapting to complex dynamic conditions, as well as open-set and 3D perception tasks.